OpenStudy (anonymous):

decide whether the following equation is a trigonometric identity. explain your reasoning. secxtanx(1-sin^2x)=sinx

myininaya (myininaya):

\[\frac{1}{cosx}\frac{sinx}{cosx}\cos^2x=sinx\] yes

OpenStudy (aravindg):

hello help me

OpenStudy (aravindg):

i hav new question

OpenStudy (anonymous):

its true because secxtanx(1-sin^2x)=sinx (tanx/cosx) cos^2x the cos cancel out so your left with tanx(cosx) tanx=sinx/cosx and you plug that for sin and you get sinx/cosx (cosx) and the cosx cancel out so your left with sinx

OpenStudy (anonymous):

thanks!